# The formation and evolution of dust in the colliding-wind binary Apep revealed by JWST

### v1

@@ Abstract
Carbon-rich Wolf-Rayet (WR) stars are significant contributors of carbonaceous dust to the galactic environment, however the mechanisms and conditions for formation and subsequent evolution of dust around these stars remain open questions. Here we present JWST observations of the WR+WR colliding-wind binary Apep which reveal an intricate series of nested concentric dust shells that are abundant in detailed substructure. The striking regularity in these substructures between successive shells suggests an exactly repeating formation mechanism combined with a highly stable outflow that maintains a consistent morphology even after reaching 0.6 pc (assuming a distance of 2.4 kpc) into the interstellar medium. The concentric dust shells show subtle deviations from spherical outflow, which could reflect orbital modulation along the eccentric binary orbit or some mild degree of non-sphericity in the stellar wind. Tracking the evolution of dust across the multi-tiered structure, we measure the dust temperature evolution that can broadly be described with an amorphous carbon composition in radiative thermal equilibrium with the central stars. The temperature profile and orbital period place new distance constraints that support Apep being at a greater distance than previously estimated, reducing the line-of-sight and sky-plane wind speed discrepancy previously thought to characterise the system.
@@
facilities: JWST(MIRI), VLT(VISIR)software: NumPy (C. R. Harris et al., 2020), SciPy (P. Virtanen et al., 2020), Matplotlib (J. D. Hunter, 2007), Astropy (Astropy Collaboration et al., 2022), JWST Science Calibration Pipeline (H. Bushouse et al., 2025)
@@ 1 Introduction
Evolved stars are significant contributors to cosmic dust. Carbon-rich Wolf-Rayet stars, the descendants of the most massive O stars ($\gtrsim 25\,\mathrm{M}_{\odot}$, P. A. Crowther, 2007), are a major producer of carbon dust in the galactic environment (R. M. Lau et al., 2020a, 2021). These stars embody the final phase of evolution before being destroyed in core-collapse supernovae. They are characterised by rapid mass loss ($\sim 10^{-5}\,\mathrm{M}_{\odot}\,\mathrm{yr}^{-1}$) driven by extreme luminosities ($\sim 10^{6}\,\mathrm{L}_{\odot}$), with a small fraction ($\lesssim$1%) of the outflow mass nucleating into dust.
@@
While significant uncertainty surrounds the exact mechanism of dust formation around WR stars, resolved imaging of dust structures around WR stars found in colliding-wind binaries have provided clues on the dust formation conditions in such systems (e.g., J. D. Monnier et al., 1999; P. G. Tuthill et al., 1999; P. Tuthill et al., 2006; R. M. Lau et al., 2020b; J. R. Callingham et al., 2019; Y. Han et al., 2020). These observations provide a link between the binary orbit, the physical properties of the stellar winds and the production of dust. For the majority of dusty systems, where the binary consists of a carbon-type WR star and a massive (O- or B-type) companion, the collision of the component stellar winds creates a dense, carbon-rich environment that favours dust nucleation (R. M. T. White & P. Tuthill, 2024). The WR component, which displays the stronger wind momentum, forces the conical shock interface to bow towards the companion. Dust forms and propagates along this shock surface, streaming radially away from the star, while at the same time orbital motion azimuthally offsets the launch direction of newer dust continuously. The net outcome of these two motions, a radially inflating wind structure injected at a rotating flow insertion point, results in a large-scale dust geometry that follows a spiral geometry.
@@
Apep (2XMM J160050.7–514245) is the most recently discovered colliding-wind WR binary to exhibit such a geometry (J. R. Callingham et al., 2019). Whereas other WR binaries typically involve an O- or B-type companion, Apep consists of two classical WR stars (J. R. Callingham et al., 2020). The WN4-6b star is thought to launch the stronger wind at 3500 km s-1 (J. R. Callingham et al., 2020), compared to 2100 km s-1 for the carbon-rich WC8 stellar wind. The combined action of these colliding winds was found to sculpt a large ($12^{\prime\prime}$) and intricate dust shell visible in the mid-infrared. Notably, a discrepancy (by a factor of $\sim$4) has been reported between the aforementioned line-of-sight wind speed measured with spectroscopy and the sky-plane expansion speed of its circumstellar dust (J. R. Callingham et al., 2019; Y. Han et al., 2020). This finding is puzzling both from the perspective of the basic kinematic physics understood for these systems, and also from studies of other colliding wind spirals (P. G. Tuthill et al., 1999, 2008; R. M. Lau et al., 2020b), both of which lead to an expectation that the proper motions of the dust expansion should be consistent with the observed spectroscopic wind speed. This has led to the suggestion that the wind velocity field of one WR component may be asymmetric, possibly due to rapid stellar rotation, which would identify Apep as a rare candidate Galactic long-duration gamma-ray burst progenitor (J. R. Callingham et al., 2019).
@@
Here we report observations of the Wolf-Rayet binary, Apep, with the James Webb Space Telescope (JWST) in the mid-infrared, targeting thermal emission from its warm ($\sim$100–300 K) circumstellar dust. This paper provides an overview of the dataset and the dynamical and thermal record of the system contained within. The observations are described in Section 2 and the data reduction in Section 3. Section 4 analyses the structure and motion of the dust structure and Section 5 models its thermal properties. Section 6 discusses these new observational constraints and their implications for estimates of the distance to Apep. The findings are summarised in Section 7. An accompanying study (R. White et al., submitted) will present a model for colliding-wind dust structures applied to Apep to constrain its orbit and the role of a tertiary massive companion in further shaping the circumstellar dust structure.
@@ 2 Observations
Instrument Filter $\lambda_{0}$ / $\Delta\lambda$ ($\mu$m) Calibrator Date
@@ 2.1 JWST/MIRI observations
We observed Apep with the MIRI Imager onboard JWST on 24 Jul 2024 UT (programme ID 5842, Y. Han et al., 2024). The observations were carried out with full aperture direct imaging using broadband filters F770W, F1500W and F2550W. The FULL detector subarray was used for all observations to maximise the field of view while the FASTR1 readout pattern was selected to minimise saturation given the choice of subarray. For each filter, one exposure was taken at each of 12 dither positions under the default CYCLING dithering pattern and LARGE pattern size, starting at set number 5 and proceeding for 3 sets. The number of groups per integration was 28, 24 and 9 for the F770W, F1500W and F2550W filters respectively, which was set by the maximum number of groups before saturation in the secondary dust shell based on its predicted surface brightness from prior modelling if such a shell were to exist (Y. Han et al., 2020). The number of integrations per exposure for the three filters was 1, 3 and 5 respectively, which was set by the total integration time required for detection of a potential secondary shell with the F770W filter and any shells close to the edge of the field of view within the F1500W and F2550W filters. An observing log is available in Table 1.
@@ 3.2 VLT/VISIR data reduction
We calibrated detector readings to flux densities (in Jy/pixel) using the accompanying observations of calibrator stars. The database of flux densities of the mid-infrared standard stars used in the calibration is maintained by ESO [^1] https://www.eso.org/sci/facilities/paranal/instruments/visir/tools.html based on M. Cohen et al. (1999).
To perform a proper motion analysis in this study, we also used archival VISIR observations from 2016 J. R. Callingham et al. (2019) and 2018 (Y. Han et al., 2020). These observations are described in the corresponding references. We re-reduced these archival datasets in the same way as the 2024 epoch to ensure consistency in subsequent analyses.
@@ 4 Dynamical evolution
Our observations of Apep detect a series of 4 spiral dust shells centred on the binary, which extend to the edge of the MIRI field of view and correspond to a separation of $1.5\times 10^{5}$ au from the binary assuming a distance of 2.4 kpc (J. R. Callingham et al., 2019, 2020; Y. Han et al., 2020). Owing to the varying wind-wind shock conditions resulting from modulation by the eccentric binary orbit (Y. Han et al., 2020; R. White et al., submitted), dust production in Apep is episodic, creating cutoff regions and segmented edges inflated to large spatial scales by the expansion of dust. The resulting rich set of well-resolved substructures within the larger nested series of shells allows us to track in detail the dust evolution between successive formation events, where all except the innermost one were previously inaccessible from ground-based infrared imaging. We describe constraints on the dynamical evolution of dust derived from such observations in this section.
@@ 4.1 Measuring dust displacement
We measured the projected radial displacement of dust shells from the central binary in the 2 VISIR filters across 3 epochs and the 3 JWST filters in 1 epoch within a series of azimuthal bins (i.e., angular sectors with the apex at the binary) that partition one full revolution of the dust shell. The angular bins were each 20 wide, and measurements were only made in bins where the outer edge of a given dust shell reflects its deprojected separation from the binary. This is expected to be the case for the majority of outermost edges within a given episode of dust (Y. Han et al., 2020; R. M. Lau et al., 2020b), as they arise from limb-brightening at the projected edges, except for the region to the north with dust absent from the otherwise azimuthally continuous distribution, possibly due to sculpting or destruction by a companion of the WR binary (R. White et al., submitted). We also avoided making measurements over a small azimuthal range to the south of the binary where the outer edge in this region is too dim for an accurate measurement in the inner shell observed with VISIR, and the southern edge of the third dust shell and beyond are outside the field of view of MIRI, preventing subsequent measurements of the expansion rate and shell spacing in this direction. The azimuthal range over which edge locations were measured are shown in Fig. 2.
Figure 2: Top: MIRI F770W image of Apep with the azimuthal range of shell radius measured in this study indicated by dashed lines, and the direction orthogonal to the line of nodes and its uncertainty range (R. White et al., submitted) indicated by dotted lines. Middle: F1500W image with points indicating examples of shell location detections. Bottom: F2550W image labelled with the E1, E2 and E3 features and apertures used to extract photometry.
@@ 4.2 Shell spacing
Given the high structural consistency across the VISIR and JWST filters, we used the inner shell’s edge location determined from the 2024 VISIR epoch across both filters, the second shell’s edge location determined from all 3 JWST filters and the third shell’s edge location determined from the F1500W and F2550W filters to measure the spacing between neighbouring dust shells in each direction. Across all 9 azimuthal bins measured, we fitted the edge location as a linear function of the shell location with a Markov chain Monte Carlo (MCMC) approach implemented with the emcee package (D. Foreman-Mackey et al., 2013), leaving the uncertainties on the edge location as a free parameter. The fits for each azimuth are displayed in Fig. 8.
@@
Figure 3: Left: Apep’s inter-shell spacing as a function of position angle. The position angle in the horizontal axis corresponds to the angle measured counterclockwise from north in Fig. 2 (left panel), in which the azimuthal range with data points are labelled with arcs. Right: the expansion speed of Apep’s inner dust shell as a function of position angle. The two curves correspond to the best-fit sinusoidal model with one and two periods over the full azimuthal range respectively. The vertical lines and shaded regions correspond to the line of nodes, apastron and the velocity vector at apastron and their uncertainties (R. White et al., submitted).
@@ 4.3 Proper motion
The inner dust shell has previously been observed by VISIR on 3 occasions between 2016 and 2018. The addition of the 2024 epoch quadruples the time baseline over which to track the expansion of the dust, which we leverage to yield significantly more reliable proper motion constraints than previous work that relied on sub-pixel edge detection methods (Y. Han et al., 2020). Note that we did not include the 2017 epoch in this analysis due to the chopping and nodding configurations used in that epoch which were not optimised for reliably imaging the extended structures in Apep.
@@ 5.1 Photometry
We selected 3 features along widely separated azimuths for which bright and consistent emission could be photometrically characterised across a wide span of shell orders and wavelengths. These features include an eastern outer arc on the bright ellipse (E1), a northern cusp on an inner edge (E2) and a western outer arc on the tail of the dust shell (E3), which are indicated in Fig. 2. We extracted the flux density of E1, E2 and E3 in shells 2 to 4 (with shell 1 being the inner shell) from all 3 JWST filters, and in shell 1 from all 3 VISIR filters across 6 images taken across 2016, 2018 and 2024. For each feature and image combination, background emission was determined using an annulus outside the outmost shell in which the feature was detected and within the same azimuthal range as the corresponding feature, and was subtracted from the flux densities measured for the given feature. Extinction correction correction was applied assuming $A_{v}=11.4$ (J. R. Callingham et al., 2019; Y. Han et al., 2020) and the extinction law determined by E. L. Fitzpatrick & D. Massa (2007). Although these apertures have been selected to minimise flux contamination, the photometric uncertainties for these specific features are still dominated by flux contributions from the faint background or foreground dust structures and the extended nebula, which we estimate to be 10%. Our calibrated flux densities for E1, E2 and E3 across the dataset are shown in Fig. 10.
@@ 5.2 Resolved SED modelling
setting $\beta=1$ for amorphous carbon dust (F. Rouleau & P. G. Martin, 1991; I. Cherchneff et al., 2000), and normalising the function with $Q_{0}=1$ and $\lambda_{0}=1$$\mu$m. Note that the specific choice of normalisation does not affect constraints on $T$, as any deviations can be absorbed by its degeneracy with the nuisance parameter, $\Omega_{\mathrm{eff}}$, during model fitting.
@@
Figure 5: The radial temperature profile of dust in Apep measured from its spatially resolved SED. The thicker error bars are derived by fitting SED models to individual features. The thinner error bars show the adjusted uncertainties upon enforcing the condition that the measurements sample the same temperature profile. The best-fit power law profile with a power-law index of -0.31 is shown, along with the blackbody equilibrium temperature, an amorphous carbon equilibrium temperature profile based on radiative transfer simulations and an amorphous carbon approximation with a power law index of -0.38 for comparison.

### v2

@@ Abstract
Carbon-rich Wolf-Rayet (WR) stars are significant contributors of carbonaceous dust to the galactic environment, however the mechanisms and conditions for formation and subsequent evolution of dust around these stars remain open questions. Here we present JWST observations of the WR+WR colliding-wind binary Apep which reveal an intricate series of nested concentric dust shells that are abundant in detailed substructure. The striking regularity in these substructures between successive shells suggests an exactly repeating formation mechanism combined with a highly stable outflow that maintains a consistent morphology even after reaching 0.6 pc (assuming a distance of 2.4 kpc) into the interstellar medium. The concentric dust shells show subtle deviations from spherical outflow, which could reflect orbital modulation along the eccentric binary orbit or non-sphericity in the stellar wind. Tracking the evolution of dust across the multi-tiered structure, we measure the dust temperature evolution that can broadly be described assuming an amorphous carbon composition in radiative thermal equilibrium with the central stars. The temperature profile and orbital period place new distance constraints that support Apep being at a greater distance than previously estimated, reducing the line-of-sight and sky-plane wind speed discrepancy previously thought to characterise the system.
@@
facilities: JWST(MIRI), VLT(VISIR)software: NumPy (numpy), SciPy (scipy), Matplotlib (matplotlib), Astropy (astropy:2022), JWST Science Calibration Pipeline (jwst_pipeline)
@@ 1 Introduction
Evolved stars are significant contributors to cosmic dust. Carbon-rich Wolf-Rayet stars, the descendants of the most massive O stars ($\gtrsim 25\,\mathrm{M}_{\odot}$, Crowther2007), are a major producer of carbon dust in the galactic environment (Lau2020a; Lau2021). These stars embody the final phase of evolution before being destroyed in core-collapse supernovae. They are characterised by rapid mass loss ($\sim 10^{-5}\,\mathrm{M}_{\odot}\,\mathrm{yr}^{-1}$) driven by extreme luminosities ($\sim 10^{6}\,\mathrm{L}_{\odot}$), with a small fraction ($\lesssim$1%) of the outflow mass nucleating into dust.
@@
While significant uncertainty surrounds the exact mechanism of dust formation around WR stars, resolved imaging of dust structures around WR stars found in colliding-wind binaries have provided clues on the dust formation conditions in such systems (e.g., Tuthill1999b; Tuthill1999; Tuthill2006b; Lau2020b; Callingham2019; Han2020). These observations provide a link between the binary orbit, the physical properties of the stellar winds and the production of dust. For the majority of dusty systems, where the binary consists of a carbon-type WR star and a massive (O- or B-type) companion, the collision of the component stellar winds creates a dense, carbon-rich environment that favours dust nucleation (White2024). The WR component, which displays the stronger wind momentum, forces the conical shock interface to bow towards the companion. Dust forms and propagates along this shock surface, streaming radially away from the star, while at the same time orbital motion azimuthally offsets the launch direction of newer dust continuously. The net outcome of these two motions, a radially inflating wind structure injected at a rotating flow insertion point, results in a large-scale dust geometry that follows a spiral geometry.
@@
Apep (2XMM J160050.7–514245) is the most recently discovered colliding-wind WR binary to exhibit such a geometry (Callingham2019). Whereas other WR binaries typically involve an O- or B-type companion, Apep consists of two classical WR stars (Callingham2020). The WN4-6b star is thought to launch the stronger wind at 3500 km s-1 (Callingham2020), compared to 2100 km s-1 for the carbon-rich WC8 stellar wind. The combined action of these colliding winds was found to sculpt a large ($12^{\prime\prime}$) and intricate dust shell visible in the mid-infrared. Notably, a discrepancy (by a factor of $\sim$4) has been reported between the aforementioned line-of-sight wind speed measured with spectroscopy and the sky-plane expansion speed of its circumstellar dust (Callingham2019; Han2020). This finding is puzzling both from the perspective of the basic kinematic physics understood for these systems, and also from studies of other colliding wind spirals (Tuthill1999; Tuthill2008; Lau2020b), both of which lead to an expectation that the proper motions of the dust expansion should be consistent with the observed spectroscopic wind speed. This has led to the suggestion that the wind velocity field of one WR component may be asymmetric, possibly due to rapid stellar rotation, which would identify Apep as a rare candidate Galactic long-duration gamma-ray burst progenitor (Callingham2019).
@@
Here we report observations of the Wolf-Rayet binary, Apep, with the James Webb Space Telescope (JWST) in the mid-infrared, targeting thermal emission from its warm ($\sim$100–300 K) circumstellar dust. This paper provides an overview of the dataset and the dynamical and thermal record of the system contained within. The observations are described in Section 2 and the data reduction in Section 3. Section 4 analyses the structure and motion of the dust structure and Section 5 models its thermal properties. Section 6 discusses these new observational constraints and their implications for estimates of the distance to Apep. The findings are summarised in Section 7. An accompanying study (White2025) will present a model for colliding-wind dust structures applied to Apep to constrain its orbit and the role of a tertiary massive companion in further shaping the circumstellar dust structure.
@@ 2 Observations
[…]
@@ 2.1 JWST/MIRI observations
We observed Apep with the MIRI Imager onboard JWST on 24 Jul 2024 UT (programme ID 5842, Han2024JWST). The observations were carried out with full aperture direct imaging using broadband filters F770W, F1500W and F2550W. The FULL detector subarray was used for all observations to maximise the field of view while the FASTR1 readout pattern was selected to minimise saturation given the choice of subarray. For each filter, one exposure was taken at each of 12 dither positions under the default CYCLING dithering pattern and LARGE pattern size, starting at set number 5 and proceeding for 3 sets. The number of groups per integration was 28, 24 and 9 for the F770W, F1500W and F2550W filters respectively, which was set by the maximum number of groups before saturation in the secondary dust shell based on its predicted surface brightness from prior modelling if such a shell were to exist (Han2020). The number of integrations per exposure for the three filters was 1, 3 and 5 respectively, which was set by the total integration time required for detection of a potential secondary shell with the F770W filter and any shells close to the edge of the field of view within the F1500W and F2550W filters. An observing log is available in Table 1.
@@ 3.2 VLT/VISIR data reduction
We calibrated detector readings to flux densities (in Jy/pixel) using the accompanying observations of calibrator stars. The database of flux densities of the mid-infrared standard stars used in the calibration is maintained by ESO [^1] https://www.eso.org/sci/facilities/paranal/instruments/visir/tools.html based on Cohen1999.
To perform a proper motion analysis in this study, we also used archival VISIR observations from 2016 Callingham2019 and 2018 (Han2020). These observations are described in the corresponding references. We re-reduced these archival datasets in the same way as the 2024 epoch to ensure consistency in subsequent analyses.
@@ 4 Dynamical evolution
Our observations of Apep detect a series of 4 spiral dust shells centred on the binary, which extend to the edge of the MIRI field of view and correspond to a separation of $1.5\times 10^{5}$ au from the binary assuming a distance of 2.4 kpc (Callingham2019; Callingham2020; Han2020). Owing to the varying wind-wind shock conditions resulting from modulation by the eccentric binary orbit (Han2020; White2025), dust production in Apep is episodic, creating cutoff regions and segmented edges inflated to large spatial scales by the expansion of dust. The resulting rich set of well-resolved substructures within the larger nested series of shells allows us to track in detail the dust evolution between successive formation events, where all except the innermost one were previously inaccessible from ground-based infrared imaging. We describe constraints on the dynamical evolution of dust derived from such observations in this section.
@@ 4.1 Measuring dust displacement
We measured the projected radial displacement of dust shells from the central binary in the 2 VISIR filters across 3 epochs and the 3 JWST filters in 1 epoch within a series of azimuthal bins (i.e., angular sectors with the apex at the binary) that partition one full revolution of the dust shell. The angular bins were each 20 wide, and measurements were only made in bins where the outer edge of a given dust shell reflects its deprojected separation from the binary. This is expected to be the case for the majority of outermost edges within a given episode of dust (Han2020; Lau2020b), as they arise from limb-brightening at the projected edges, except for the region to the north with dust absent from the otherwise azimuthally continuous distribution, possibly due to sculpting or destruction by a companion of the WR binary (White2025). We also avoided making measurements over a small azimuthal range to the south of the binary where the outer edge in this region is too dim for an accurate measurement in the inner shell observed with VISIR, and the southern edge of the third dust shell and beyond are outside the field of view of MIRI, preventing subsequent measurements of the expansion rate and shell spacing in this direction. The azimuthal range over which edge locations were measured are shown in Fig. 2.
Figure 2: Top: MIRI F770W image of Apep with the azimuthal range of shell radius measured in this study indicated by dashed lines, and the direction orthogonal to the line of nodes and its uncertainty range (White2025) indicated by dotted lines. Middle: F1500W image with points indicating examples of shell location detections. Bottom: F2550W image labelled with the E1, E2 and E3 features and apertures used to extract photometry.
@@ 4.2 Shell spacing
Given the high structural consistency across the VISIR and JWST filters, we used the inner shell’s edge location determined from the 2024 VISIR epoch across both filters, the second shell’s edge location determined from all 3 JWST filters and the third shell’s edge location determined from the F1500W and F2550W filters to measure the spacing between neighbouring dust shells in each direction. Across all 9 azimuthal bins measured, we fitted the edge location as a linear function of the shell location with a Markov chain Monte Carlo (MCMC) approach implemented with the emcee package (emcee), leaving the uncertainties on the edge location as a free parameter. The fits for each azimuth are displayed in Fig. 8.
@@
Figure 3: Left: Apep’s inter-shell spacing as a function of position angle. The position angle in the horizontal axis corresponds to the angle measured counterclockwise from north in Fig. 2 (left panel), in which the azimuthal range with data points are labelled with arcs. Right: the expansion speed of Apep’s inner dust shell as a function of position angle. The two curves correspond to the best-fit sinusoidal model with one and two periods over the full azimuthal range respectively. The vertical lines and shaded regions correspond to the line of nodes, apastron and the velocity vector at apastron and their uncertainties (White2025).
@@ 4.3 Proper motion
The inner dust shell has previously been observed by VISIR on 3 occasions between 2016 and 2018. The addition of the 2024 epoch quadruples the time baseline over which to track the expansion of the dust, which we leverage to yield significantly more reliable proper motion constraints than previous work that relied on sub-pixel edge detection methods (Han2020). Note that we did not include the 2017 epoch in this analysis due to the chopping and nodding configurations used in that epoch which were not optimised for reliably imaging the extended structures in Apep.
@@ 5.1 Photometry
We selected 3 features along widely separated azimuths for which bright and consistent emission could be photometrically characterised across a wide span of shell orders and wavelengths. These features include an eastern outer arc on the bright ellipse (E1), a northern cusp on an inner edge (E2) and a western outer arc on the tail of the dust shell (E3), which are indicated in Fig. 2. We extracted the flux density of E1, E2 and E3 in shells 2 to 4 (with shell 1 being the inner shell) from all 3 JWST filters, and in shell 1 from all 3 VISIR filters across 6 images taken across 2016, 2018 and 2024. For each feature and image combination, background emission was determined using an annulus outside the outmost shell in which the feature was detected and within the same azimuthal range as the corresponding feature, and was subtracted from the flux densities measured for the given feature. Extinction correction correction was applied assuming $A_{v}=11.4$ (Callingham2019; Han2020) and the extinction law determined by Fitzpatrick2007. Although these apertures have been selected to minimise flux contamination, the photometric uncertainties for these specific features are still dominated by flux contributions from the faint background or foreground dust structures and the extended nebula, which we estimate to be 10%. Our calibrated flux densities for E1, E2 and E3 across the dataset are shown in Fig. 10.
@@ 5.2 Resolved SED modelling
[…]
@@
Figure 5: The radial temperature profile of dust in Apep measured from its spatially resolved SED. The thicker error bars are derived by fitting SED models to individual features. The thinner error bars show the adjusted uncertainties upon enforcing the condition that the measurements sample the same temperature profile. The best-fit power law profile with a power-law index of -0.31 is shown, along with the blackbody equilibrium temperature, an amorphous carbon equilibrium temperature profile based on RADMC-3D simulations and an amorphous carbon approximation with a power law index of -0.38 for comparison.